# Polycyclic codes over serial rings and their annihilator CSS construction

### v1

@@ 8 Conclusion
In this paper, we have investigated the structure of $f(x)$-polycyclic codes over a specific category of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial. We have established a unified framework in the literature of constacyclic and polycyclic codes over various rings, all of which are examples of the ring $\mathscr{R}$. We have defined the new concepts QsDP codes and $f(x)$-polycyclic-QsDP codes, and have proved that $f(x)$-polycyclic codes over $\mathscr{R}$ are isomorphic to $f(x)$-polycyclic-QsDP codes. Additionally, we have established a polynomial structure for the Eclidean dual of $f(x)$-polycyclic codes over $\mathscr{R}$. The annihilator dual is an alternative duality concept ensuring that the dual of $f(x)$-polycyclic codes has an $f(x)$- polycyclic structure, a feature that fails in Euclidean dual. We have studied some properties of the annihilator dual of $f(x)$-polycyclic code over the ring $\mathscr{R}$ and have provided necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator LCD, and annihilator dual-containing polycyclic codes over $\mathscr{R}$. Finally, we have provided an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes generated by a single element $g(x)$, where $g(x)$ divides $f(x)$.

### v3

@@ 8 Conclusion and future work
In this paper, we have investigated the structure of $f(x)$-polycyclic codes over a specific category of serial rings, defined as $\mathscr{R}=R[x_{1},\ldots,x_{s}]/\langle t_{1}(x_{1}),\ldots,t_{s}(x_{s})\rangle$, where $R$ is a chain ring and each $t_{i}(x_{i})$ in $R[x_{i}]$ is a monic square-free polynomial. We have established a unified framework in the literature of constacyclic and polycyclic codes over various rings, all of which are examples of the ring $\mathscr{R}$. We have defined the new concepts QsDP codes and $f(x)$-polycyclic-QsDP codes, and have proved that $f(x)$-polycyclic codes over $\mathscr{R}$ are isomorphic to $f(x)$-polycyclic-QsDP codes. Additionally, we have established a polynomial structure for the Euclidean dual of $f(x)$-polycyclic codes over $\mathscr{R}$. The annihilator dual is an alternative duality concept ensuring that the dual of $f(x)$-polycyclic codes have an $f(x)$- polycyclic structure, a feature that fails in the Euclidean dual. We have studied some properties of the annihilator dual of $f(x)$-polycyclic code over the ring $\mathscr{R}$ and have provided necessary and sufficient conditions for the existence of annihilator self-dual, annihilator self-orthogonal, annihilator LCD, and annihilator dual-containing polycyclic codes over $\mathscr{R}$. Finally, we have provided an annihilator CSS construction to derive quantum codes from annihilator dual-preserving $f(x)$-polycyclic codes over Frobenius rings. The use of a general polynomial other than $x^{n}-1$ or $x^{n}-\lambda$ (cyclic and constacyclic case) provides the chance to explore a broader range of factors, and the annihilator duality provides that the dual code is always in the same ambient space and it is easy to compute, hence the annihilator CSS construction provides a broader class of codes where to find quantum codes. As a future work, it will be worth studying the annihilator duality and checking if it yields new quantum codes. Also in [7], the authors have recently studied quasi-polycyclic codes and their application for constructing quantum error-correcting codes based on the CSS construction for Euclidean and symplectic duality. Extending these results to the annihilator dual setting may yield new classes of quantum codes.